OpenStudy (df001):

(Algebra) What to do when index is 6 and radical 1,000,000?

OpenStudy (df001):

\[\sqrt[6]{1,000,000}\]

OpenStudy (df001):

can someone show me how to solve this in steps pls I only know how to do it when the index is 2

OpenStudy (df001):

I need help pls

OpenStudy (df001):

hi i need help pls

OpenStudy (df001):



OpenStudy (anonymous):

What are the factors of 1,000,000?

OpenStudy (df001):

1000 x 1000

OpenStudy (df001):

i only know how to do when index is 2

OpenStudy (anonymous):

\[\sqrt[6]{1000000}=\sqrt[6]{^{10^6}}=10\]

OpenStudy (df001):

but i never done one with a 6

OpenStudy (df001):

I know the answer is 10 but, im not sure how to get it

OpenStudy (anonymous):

@amirreza1870 Giving out direct answers is against the Code of Conduct of openstudy. :)

OpenStudy (df001):

i dont know how to reduce it